Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (T. Venturi at Sun Jun 7 09:43:49 METDST 1998) Success
Experiment successful. No apparent problems reported.

Onsala (John Conway at Sun Jun 7 17:58:41 METDST 1998) Success
No known problems

Torun (K. Borkowski at Mon Jun 8 09:26:07 METDST 1998) Success
No known problems.

Westerbork (Tony Foley at Mon Jun 8 10:05:18 METDST 1998) Minor failures
Missed a scan of 1328_307 at 02:41 as it took too long to slew to it. Otherwise looks OK.

Noto (C.Trigilio at Mon Jun 8 18:07:27 METDST 1998) Success

Jodrell (Alastair Gunn at Tue Jun 9 03:07:37 METDST 1998) Success
Strong RFI in VCs 5 and 6. FS clock a few seonds slow. No other known problems.

Cambridge (Alastair Gunn at Tue Jun 9 03:08:03 METDST 1998) Minor failures
Only partial signal in BBCs 3 & 4 due to microwave link limitations. Vacuum failed at 04:56 UT, restarted for scan at 05:29 UT.

Calibration (Jean\-Francois Desmurs at Thu Jun 11 12:37:08 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the mai98 directory under the name gp017.antab. Rem: Jb strange Tsys values. Cm No Tsys. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 14:58:16 METDST 1998) Minor failures
Should have been scheduled for VLBA terminal. Mk4 write quality on thin tape poor, particularly hi-numbered tracks.